can you help me find the correct answer.
find the center and the radius of the circle
x^2+y^2+6y-27=0

Answer by stanbon(48568)

find the center and the radius of the circle
x^2+y^2+6y-27=0
-------------
Complete the square on the x-terms and the y-terms respectively:
---------
[x^2] + [y^2+6y + ?] = 27+?
(x-0)^2 + (y^2+6y+(6/2)^2) = 27 + 9
(x-0)^2 + (y+3)^2 = 36
Center at (0,-3)
Radius= 6
--------------
cheers,
Stan H.